Crane Corporation has one temporary difference at the end of 2020 that will reverse and cause taxable amounts of $56,200 in 2021, $61,000 in 2022, and $66,500 in 2023. Crane’s pretax financial income for 2020 is $285,200, and the tax rate is 30% for all years. There are no deferred taxes at the beginning of 2020.

Compute taxable income and income taxes payable for 2020.

Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2020.

Prepare the income tax expense section of the income statement for 2020, beginning with the line “Income before income taxes.”
Answer #1

a)Total reversal amount = 56200+61000 +66500 = 183700

Taxable income =Pretax financial income -Future taxable amount

                     = 285200-183700

                      = 101500

Income taxes payable =Taxable income *tax rate

                      = 101500*30%

                    = 30450

b)

Date Account title Debit credit
December 31 ,2020 Income tax expense (285200*30%) 85560
Deferred tax liability 55110
Income tax payable 30450

#Out of deferred tax liability ,56200*30%= 16860 will be classified as current (as same is payable within one year from reporting period) and rest will be classified as non current :55110-16860= 38250

c)

Income before income tax 285200
less:Income tax expense (85560)
Net income 199640